OpenStudy (anonymous):

The sum of two numbers is 72, and twice their difference is 24. What is the smaller of the two numbers?

OpenStudy (anonymous):

suppose the two numbers are x and y. what can you tell about x and y from the description?

OpenStudy (anonymous):

x + y= 72

OpenStudy (anonymous):

okay and?

OpenStudy (anonymous):

2(x-y)=24

OpenStudy (anonymous):

This is where I get stuck

OpenStudy (anonymous):

okay good. so x+y = 72 and 2(x-y)=24 or x-y = 12 so you have x+y = 72 x-y =12 solve for x and y

OpenStudy (anonymous):

I'm not sure where to go from there. Do I combine the two equations?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

well lets see, what could we get for x, that will make 72 if we add y? Plug n chug works, lets try 40 for X.

OpenStudy (anonymous):

What can we add to 40 to make 72? 32.

OpenStudy (anonymous):

42+30=72

OpenStudy (anonymous):

42-30=12, 12x2=24

OpenStudy (anonymous):

made a mistake, i took 40 instead of 40, but if we were to pick 40, we would still see that it wont meet the requirements, so that is wrong anyway.

OpenStudy (anonymous):

blah, instead of 42**

OpenStudy (anonymous):

ok. I understand it a little bit more

OpenStudy (anonymous):

brittany, you had x+y = 72 x - y = 12 add the two equations 2x = 84. so x = 42. if x = 42, then 42+y = 72. so y = 30. so your answer is 30

OpenStudy (anonymous):

I concur.

OpenStudy (anonymous):

Alright, thanks guys

OpenStudy (anonymous):

:)